# Test-Time Degradation Adaption for Open-Set Image Restoration

### v3

@@ Abstract
In contrast to close-set scenarios that restore images from a predefined set of degradations, open-set image restoration aims to handle the unknown degradations that were unforeseen during the pretraining phase, which is less-touched as far as we know. In this work, we explicitly study this challenging problem and reveal its essence, i.e., the unidentified distribution shifts between test and training data. In recent, test-time adaptation emerges as a fundamental method to address this inherent disparities. Inspired by this, we propose a test-time degradation adaption framework for open-set image restoration, which involves three components, i.e., i) a pre-trained and degradation-agnostic diffusion model to generate clean images, ii) a test-time degradation adapter adapts the unknown degradations based on the input image during the testing phase, and iii) the adapter-guided image restoration guides the model through the adapter to produce the corresponding clean image. Through experiments on multiple degradations absent from the training data, we show that our method achieves comparable even better performance than those task-specific methods.
@@ 1 Introduction
In recent years, significant advances have been made in the realm of image restoration, and demonstrated remarkable capabilities in addressing a multitude of image degradations [9, 17, 12, 44, 24]. However, a common limitation among most existing methods is their reliance on assumptions rooted in a close-set scenario, i.e., the test degradations closely resemble those encountered during the pretraining phase. This assumption restricts the applicability of these methods to specific types of degradation, hindering their adaptability to a broader range of real-world scenarios where diverse and unforeseen degradations may arise.
@@
Figure 1: The differences of image restoration (IR) tasks. To be specific, (a) Classic IR works in a close-set scenario where the training and test degradations are the same and known, and customizes a specialized model for each one. (b) All-in-one IR also works in a close-set scenario where the training and test degradations are the same but unknown, and addresses them through a single model. (c) Zero-shot IR focuses on recovery from single degraded sample, which is free from the training degradations, but often suffers from the priors of the test degradations. In contrast, (d) OIR works in a open-set scenario where the test degradations are unknown and different from the pretraining ones. This is analogous to the problem in the realm of natural language processing, which applies the pre-trained Large Language Model to the various downstream tasks not predefined during the pretraining phase.
@@
To break through the limitation, this study delves into a more challenging and less-touched problem, i.e., open-set image restoration (OIR). In contrast to conventional close-set approaches, where models only address specific degradations encountered during the pretraining phase, OIR poses a formidable challenge by requiring models to handle unknown degradations absent from the training data. This shift in focus aims to push the boundaries of image restoration, fostering the development of models that exhibit adaptability and resilience in the face of diverse and unexpected degradations. By tackling OIR problem, this work seeks to pave the way for more robust and versatile solutions applicable to complex degradations in the real-world scenarios.
@@
The essence of OIR lies in addressing the challenges posed by real-world scenarios, where the test data may differ significantly from the training data. This could involve variations in degradation types and levels, lighting conditions and so on that were not explicitly covered during the pretraining phase. To tackle this challenge, test-time adaption (TTA) recently emerges as an effective methodology to address the inherent disparities between test and training data [37]. It adapts the pre-trained model during the testing phase based on the specific characteristics of the test data, allowing the model to perform better on a wider range of input scenarios. This is particularly useful in OIR where the test data may vary from the training data, and the pre-trained model needs to adapt to unknown and unseen degradations for optimal performance.
@@
Building on this inspiration, we present a Test-time degradation Adaption framework for Open-set image restoration, termed TAO, which harnesses the idea of TTA to provide a robust solution for OIR. Specifically, TAO consists of three components, i.e., a pre-trained image diffusion model (PDM), a test-time degradation adapter (TDA), and the adapter-guided image restoration (AIR). In this framework, PDM is adopted as the foundation model for OIR due to the following considerations. First, PDM captures rich knowledges of generating various high-quality visual scenarios, which could be regarded as a generic pretraining for OIR targeting at producing clean images. Second, PDM is degradation-agnostic, and any degradations in test data could be considered as unforeseen. Adapting PDM for OIR during the testing phase has a great potential to pave the way towards universal image restoration system. Note that the testing phase of PDM is referred to the process of image generation, i.e., the step-by-step denoising process.
@@
After each denoising step, TDA and AIR are sequentially conducted for adapting to open-set scenarios and guiding image restoration, respectively. Specifically, TDA employs a learnable adapter during the testing phase for adapting PDM on the test degraded image. This adapter is devised for domain alignment that aligns the generative domain of PDM to the degraded domain of the test image. In this way, the generated image could be translated to the corresponding degraded one, which could be further supervised toward the test degraded image, and in turn updating the generated image. AIR is elaborated to conduct this supervised updating for image restoration. In brief, AIR is inspired by the observation that PDM exhibits an intriguing temporal dynamic [5] during the step-by-step denoising process. Therefore, AIR dynamically and accordingly adjusts the supervision strategies at different denoising steps for achieving a better image restoration.
Overall, the contributions of this work could be summarized as follows.
@@
  • To the best of our knowledge, this could be one of the first work that explicitly studies OIR and discovers its essence, i.e., the unidentified distribution shifts between test and training data.
@@
  • This work reveals that TTA is an effective methodology for OIR, by adapting the pre-trained model on the test data during the test phase, to address the inherent disparities between test and training data.
@@ 2 Related Work
In this section, we briefly review recent advances in related topics, which mainly involves test-time adaption, all-in-one image restoration, and zero-shot image restoration.
@@
Test-time Adaption has shown to be effective at tackling distribution shifts between test and training data, by adapting the pre-trained model on test samples. In the past several years, it has gained increasing attention and plentiful methods have been proposed. For example, [37, 42] updated the specific model parameters through the test samples by resorting to the unsupervised objectives. Similarly, [38, 10] aimed to solve the continually changing distribution shifts along the test time. Moreover, [30, 45] considered more challenging and practical adaption settings such as single sample, label shifts, and mixed domain shifts. Besides, [33, 20] shifted their focuses on applications beyond image recognition, such as semantic segmentation and pose estimation.
@@
This work focuses on application of solving OIR through TTA methodologies. In brief, we incorporate PDM with an adapter which is updated for adapting to the test degraded images during the test phase. Besides, we also consider the challenging adaption setting of single test sample.
@@
All-in-one Image Restoration is an emerging problem which has been attracted more and more attentions. Specifically, it aims to address the unknown degradations within a predefined set through one model. For example, [25] firstly studied this problem and proposed to automatically extract the degradation representation from the unknowingly degraded images for assisting the restoration. [31] aimed to encode the degradations’ information into prompts, which are then used to dynamically guide the restoration. [17] focused on using the generated text prompts to guide the latent diffusion process for restoring images with unknown degradations. Although remarkable progress has been achieved, they usually need to be trained on a predefined set of degradations, and recover images from the degradations that are unknown but within this set. In other words, they could only be applied to the close-set scenarios.
@@
In contrast, this work is devoted to address OIR problem, where the test degradations are unknown during the test phase, and meanwhile unforeseen during the pretraining phase. Note that OIR is established on a pre-trained model for image restoration, and adapts it to the downstream tasks (i.e., unknown and unforeseen degradations) during the test phase. Here, we empirically regard the diffusion model [15] as a generic pretraining task for various image restoration tasks.
@@
Zero-shot Image Restoration focuses on recovering the clean image from a single degraded one, without relying on paired clean-degraded images for training. Classic methods are usually devised for specific tasks. For example, [34] performed super resolution by downsampling the low-resolution image to construct the lower-low training pairs. [23] trained dehazing networks by decoupling and coupling the haze image through the physical model of haze formation. Recently, some works focused on methods with more general purpose. For instance, [18, 39] applied the decomposition approaches on the predefined linear degradations, and used the diffusion model to address multiple tasks of image restoration. Similarly, [9] introduced the conditional guidance for eliminating linear or blind degradations through predefined or optimizable linear degradation operators, respectively.
@@
This work is different from the above methods on both the problem and solution. On problem, zero-shot focuses on the restoration from single degraded sample, while open-set focuses on handling the unknown degradations that were unforeseen during the pretraining phase, where the degraded sample does not have to be a single one. On solution, the above methods usually exploit additional degradation priors to assist restoration from single degraded image, while our method only introduces a degradation-agnostic adapter for all unknown and unforeseen degradations.
@@ 3 The Proposed Method
In this section, we first briefly retrospect the preliminary knowledge of PDM, and then progressively illustrate the proposed TAO, TDA and AIR.
@@ 3.1 Preliminary
Given a data distribution $x_{0}\sim q(x_{0})$ and a noise distribution $x_{T}\sim\mathcal{N}(0,I)$, the diffusion model [15] defines a $T$-timestep diffusion process, which corrupts $x_{0}$ to $x_{T}$ by sequentially adding the random noise $\epsilon\sim\mathcal{N}(0,I)$, and a $T$-timestep denoising process, which recovers $x_{T}$ to $x_{0}$ by progressively removing the noise $\epsilon$. Importantly, there is an elegant property in the diffusion process, i.e.,
@@
where $t\in\{0,...,T\}$, $\bar{\alpha}_{t}=\prod^{t}_{i=0}\alpha_{i}$, $\alpha_{i}=1-\beta_{i}$ and $\beta_{i}$ is the variance of the $i$th timestep. Through this formula, the diffusion model learns a denoiser for the denoising process via the following objectives, i.e.,
@@
where $\mathcal{Z}_{\theta}$ is the denoising network, and $t$ is randomly sampled during the training phase. For a well-trained diffusion model, denoising process progressively yields $x_{t-1}$ from $x_{t}$ in terms of $\epsilon_{t}=\mathcal{Z}_{\theta}(x_{t},t)$ until $t=0$. According to Bayes Theorem, $x_{t-1}$ could be sampled through the following process, i.e.,
@@
where $\hat{x}_{0}$ is the estimated output of $x_{t}$ in terms of Eq.(1).To guide $x_{t-1}$ towards the desired output $x_{0}$, the conditional guidance $y$ could be introduced through the following formula [8, 9], i.e.,
@@
3.2 The Framework of TAO
Figure 2: Overview of the proposed TAO, which addresses OIR by exploiting i) a PDM as the generic pre-trained model for various tasks of image restoration, and at each step of denoising alternatively performing ii) a TDA for adapting to the unknown and unseen degradations posed by open-set scenario, and iii) an AIR for guiding the PDM through the TDA to restore the desired clean images.
To endow PDM with the capacity of OIR, the key lies in implementing $p_{\phi}(\cdot)$ to produce appropriate guidance of gradients $\mathcal{G}$, towards the clean version $x_{0}$ of the test degraded image $y$. In this work, $p_{\phi}(\cdot)$ is implemented through two components, i.e., TDA and AIR, which are devised for open-set scenarios and image restoration, respectively. To address the challenges posed by open-set scenarios, TDA introduces a degradation-agnostic adapter during the test phase (i.e., denoising process of PDM) for adapting PDM to the unknown and unforeseen degradations. The adapter is implemented by a simple neural network, and optimized once at each denoising timestep towards aligning the domain of generated images to that of the test degraded image. As a result, the generated images could be translated to the corresponding degraded ones, which could be further supervised towards the $y$, and in turn produce $\mathcal{G}$ to update the generated images towards the desired $x_{0}$. AIR is elaborated to conduct this supervised process for image restoration. Specifically, AIR is inspired by the observation that PDM exhibits an intriguing temporal dynamic during the denoising process, i.e., images are generated from the unrecognizable contents to the perceptually rich contents, and finally the imperceptible details. Therefore, AIR dynamically adjusts the supervision strategies by gradually shifting the focuses from high- to low-level contents during the denoising process.
The framework of TAO is illustrated in Fig.2, where an PDM is frozen to progressively denoising the pure noise $x_{T}$ to the clean image $x_{0}$. For a given degraded image $y$, TAO first samples two maps of random noise at $t=T$, and thus two images will be generated at each timestep. Here, we refer to the image guided towards the clean version of the $y$ as the guided image $x^{g}_{t}$, and the other unguided one as the referred image $x^{r}_{t}$. The introduction of $x^{r}_{t}$ is for preventing TDA from falling into trivial solutions, and assisting AIR to guide $x^{g}_{t}$ towards the desired clean image $x^{g}_{0}$. At each timestep $t$, the images $x^{g}_{t}$, $x^{r}_{t}$, and $y$ are first fed into TDA for optimizing the adapter $\phi$ through the objectives of domain alignment, and then sent into AIR for updating $x^{g}_{t}$ towards $x^{g}_{0}$, which best matches the $y$ undergoing the $\phi$, through the conditional guidance in Eq.(4).

### v4

# Test-Time Degradation Adaptation for Open-Set Image Restoration

@@ Abstract
In contrast to close-set scenarios that restore images from a predefined set of degradations, open-set image restoration aims to handle the unknown degradations that were unforeseen during the pretraining phase, which is less-touched as far as we know. This work study this challenging problem and reveal its essence as unidentified distribution shifts between the test and training data. Recently, test-time adaptation has emerged as a fundamental method to address this inherent disparities. Inspired by it, we propose a test-time degradation adaptation framework for open-set image restoration, which consists of three components, i.e., i) a pre-trained and degradation-agnostic diffusion model for generating clean images, ii) a test-time degradation adapter adapts the unknown degradations based on the input image during the testing phase, and iii) the adapter-guided image restoration guides the model through the adapter to produce the corresponding clean image. Through experiments on multiple degradations, we show that our method achieves comparable even better performance than those task-specific methods. The code is available at https://github.com/XLearning-SCU/2024-ICML-TAO.
Image Restoration, Test-Time Adaptation, Diffusion Model
@@ 1 Introduction
In recent years, significant advances have been made in the realm of image restoration, and have demonstrated remarkable capabilities in addressing a multitude of image degradations (Fei et al., 2023; Jiang et al., 2023; Gou et al., 2022; Zhao et al., 2023; Li et al., 2021b). However, a common limitation among most existing methods is their reliance on assumptions rooted in a close-set scenario, i.e., the test degradations closely resemble those encountered during the pretraining phase. This assumption restricts the applicability of these methods to specific types of degradation, hindering their adaptability to a broader range of real-world scenarios where diverse and unforeseen degradations may arise.
@@
Figure 1: The differences of image restoration (IR) tasks. To be specific, (a) Classic IR works in a close-set scenario where the training and test degradations are the same and known, and customizes a specialized model for each one. (b) All-in-one IR also works in a close-set scenario where the training and test degradations are the same but unknown, and addresses them through a unified model. (c) Zero-shot IR focuses on recovery from single degraded image, which is free from the training degradations, but often requires priors about the test degradations in advance. In contrast, (d) OIR works in an open-set scenario where the test degradations are unknown and different from the pretraining ones. This is analogous to the challenge in natural language processing, which applies the pre-trained large language model to the various downstream tasks not predefined during the pretraining phase.
@@
To break the above assumption, this work delves into the more challenging and less-touched problem of open-set image restoration (OIR). In contrast to conventional close-set approaches, where models only address the specific degradations encountered during the pretraining phase, OIR poses a formidable challenge by requiring models to handle the unknown degradations absent from the training data. This shift in focus aims to push the boundaries of image restoration, fostering the development of models that exhibit adaptability and resilience in the face of diverse and unexpected degradations. By tackling the OIR problem, this work seeks to pave the way for more robust and versatile solutions applicable to complex degradations in the real-world scenarios.
@@
The essence of OIR lies in addressing the challenges posed by real-world scenarios, where the test data may differ significantly from the training data. This could involve variations in degradation types and levels, lighting conditions and so on that were not explicitly covered in the training data. To tackle this challenge, test-time adaptation (TTA) has recently emerged as an effective methodology to address the inherent disparities between the test and training data (Wang et al., 2021). Specifically, it adapts the pre-trained model during the testing phase based on the specific characteristics of the test data, allowing the model to perform better on a wider range of input scenarios. This is particularly useful in OIR where the test data may vary from the training data, and the pre-trained model should to adapt to unknown and unseen degradations for optimal performance.
@@
Building on this inspiration, we present a Test-time degradation Adaptation framework for Open-set image restoration, dubbed TAO, which harnesses the idea of TTA to provide a robust solution for OIR. Specifically, TAO consists of three components, i.e., a pre-trained image diffusion model (PDM), a test-time degradation adapter (TDA), and the adapter-guided image restoration (AIR). In this framework, PDM is adopted as the foundation model for OIR due to the following considerations. First, PDM captures rich knowledges of generating various high-quality visual scenarios, which could be regarded as a generic pretraining for OIR targeting at producing clean images. Second, PDM is degradation-agnostic and any degradations in the test data could be considered as unforeseen. Note that the test phase of PDM is referred to the process of image generation, i.e., the step-by-step denoising process.
@@
After each denoising step, TDA and AIR are sequentially conducted for adapting to open-set scenarios and guiding image restoration, respectively. Specifically, TDA employs a learnable adapter during the test phase for adapting PDM on the test degraded image. This adapter is devised for domain alignment that aligns the generative domain of PDM to the degraded domain of the test image. In this way, the generated clean image could be translated to the corresponding degraded one, which could be further supervised towards the test degraded image, and in turn updating the generated clean image. AIR is elaborated to conduct this supervised updating for image restoration. In brief, AIR is inspired by the observation that PDM exhibits an intriguing temporal dynamic (Choi et al., 2022) during the step-by-step denoising process. Therefore, AIR dynamically and accordingly adjusts the supervision strategies at different denoising steps for achieving a better performance.
Overall, our contributions could be summarized as follows.
@@
  • As far as we know, this could be one of the first work that explicitly studies OIR and discovers its essence, i.e., the unidentified distribution shifts between the test and training data.
@@
  • This work reveals that TTA is an effective methodology for OIR, by adapting the pre-trained model on the test data during the test phase, to address the inherent disparities between the test and training data.
@@ 2 Related Work
In this section, we briefly review recent advances in related topics, which mainly involves test-time adaptation, all-in-one image restoration, and zero-shot image restoration.
@@
Test-time adaptation has shown to be effective at tackling distribution shifts between test and training data, by adapting the pre-trained model on test samples. In the past several years, it has gained increasing attentions and plentiful methods have been proposed. For example, (Wang et al., 2021; Yang et al., 2024) updated the specific model parameters through the test samples by resorting to the unsupervised objectives. Similarly, (Wang et al., 2022a; Gan et al., 2023) aimed to solve the continually changing distribution shifts along the test time. Moreover, (Niu et al., 2023; Zhou et al., 2023) considered more challenging and practical adaptation settings such as single sample, label shifts, and mixed domain shifts. In addition, (Shin et al., 2022; Lee et al., 2023) shifted their focuses on applications beyond image recognition, such as semantic segmentation and pose estimation.
@@
This work focuses on application of solving OIR through TTA methodologies. In brief, we incorporate PDM with an adapter which is updated for adapting to the test degraded image during the test phase. Besides, we also consider the challenging adaptation setting of single test sample.
@@
All-in-one Image Restoration is an emerging issue which has been attracted more and more attentions. Specifically, it aims to address the unknown degradations within a predefined set through one model. For example, (Li et al., 2022) firstly studied this problem and proposed to automatically extract the degradation representation from the unknowingly degraded images for assisting their restoration. (Potlapalli et al., 2023) aimed to encode the degradation information into prompts which are then used to dynamically guide the restoration. (Jiang et al., 2023) focused on using the generated text prompts to guide the latent diffusion model for restoring images with unknown degradations. Although remarkable progress has been achieved, they usually need to be trained on a predefined set of degradations, and recover images from the degradations that are unknown but within this set. In other words, they could only be applied to the close-set scenarios.
@@
In contrast, this work is devoted to address OIR challenge, wherein the test degradations are unknown during the test phase, meanwhile unforeseen during the pretraining phase. Note that OIR is established on the pre-trained model for image restoration, and adapts the model to the downstream tasks (i.e., unknown and unforeseen degradations) during the test phase. This work empirically regards the diffusion model (Ho et al., 2020) as a generic pretraining task for various image restoration tasks.
@@
Zero-shot Image Restoration focuses on recovering the clean image from a single degraded one, without relying on paired clean-degraded images for training. Classic methods are usually devised for specific tasks. For example, (Shocher et al., 2018) performed super resolution by downsampling the low-resolution image to construct the lower-low training pairs. (Li et al., 2021a) trained dehazing networks by decoupling and coupling the haze image through the physical model of haze formation. In recent, some works focused on methods with more general purpose. For instance, (Kawar et al., 2022; Wang et al., 2022c) applied the decomposition approaches on the predefined linear degradations, and used the diffusion model to address multiple tasks of image restoration. Similarly, (Fei et al., 2023) introduced a conditional guidance for eliminating linear and blind degradations through the predefined and optimizable linear degradation operators, respectively.
@@
This work is different from the above methods on both the problem and solution. On the problem, zero-shot focuses on the restoration from single degraded sample, while open-set focuses on dealing with the unknown degradations that were unforeseen during the pretraining phase, where the degraded sample does not have to be a single one. On the solution, the above methods usually exploit additional degradation priors to assist restoration from single degraded image, while our method only introduces a degradation-agnostic adapter for all unknown and unforeseen degradations.
@@ 3 The Proposed Method
In this section, we first briefly review the preliminary knowledges, and then progressively illustrate the proposed TAO and its two core components of TDA and AIR.
Figure 2: Overview of the proposed method, which exploits (i) a PDM as the generic pre-trained model for OIR. After each denoising step, it first performs (ii) TDA for adapting to the unknown and unseen degradations posed by open-set scenarios, and then conducts (iii) AIR for optimizing the guided image towards the restored clean image. Note: the snowflake icon indicates the image or model is fixed, and the flame icon indicates the image or model will be updated through the gradients.
@@ 3.1 Preliminary
Open-set image restoration is devoted to handle the unknown degradations that were also unforeseen during the pretraining phase. Mathematically, for a given model $M$ pretrained on the distribution $D_{train}$, OIR requires to restore a test degraded image $y\sim D_{test}$ so that $L(M(y),x)$ is minimal, wherein $D_{test}$ is unknown and different from $D_{train}$, $L$ is a function to measure the discrepancy, and $x$ is the clean version of $y$. Formally, OIR essentially poses a problem of the unidentified distribution shifts between the test and training data. For OIR, two fundamental principles should be followed. (i) Unknown: OIR is agnostic to degradations during the test phase, and any prior knowledge about specific degradations is prohibitive. ii) Unforeseen: OIR is independent of a specific pretraining task, and only requires the test degradations were not presented for pretraining.
Test-time adaptation aims to improve the model’s performance on new, unseen data by adjusting its parameters based on the test samples. Mathematically, for a given model $M$ pretrained on the distribution $D_{train}$, TTA optimizes some parameters $\theta$ from $M$ on the test samples $y\sim D_{test}$ through the objective function $\min_{\theta}L(y;\theta)$, wherein $D_{test}$ is different from $D_{train}$. For TTA, two fundamental principles should be followed. (i) TTA starts with a pre-trained model which has learned general knowledges for the target task. (ii) TTA works on the test data where the training pairs are unavailable. In this work, we adopted the diffusion model as the pre-trained model, the generated and degraded images as the training pairs to implement TTA for OIR task.
The Diffusion Model. Given a data distribution $x_{0}\sim q(x_{0})$ and a noise distribution $x_{T}\sim\mathcal{N}(0,I)$, diffusion model (Ho et al., 2020) defines a $T$-timestep diffusion process, which corrupts $x_{0}$ to $x_{T}$ by sequentially adding the random noise $\epsilon\sim\mathcal{N}(0,I)$, and a $T$-timestep denoising process, which recovers $x_{T}$ to $x_{0}$ by progressively eliminating the noise $\epsilon$. Importantly, there is an elegant property in the diffusion process, i.e.,
@@
where $t\in\{0,...,T\}$, $\bar{\alpha}_{t}=\prod^{t}_{i=0}\alpha_{i}$, $\alpha_{i}=1-\beta_{i}$ and $\beta_{i}$ is the variance of the $i$th timestep. Based on this formula, the diffusion model learns a denoiser for the denoising process via the following objective, i.e.,
@@
where $\mathcal{Z}_{\theta}$ is the denoising network, and $t$ is randomly sampled during the training phase. For a well-trained diffusion model, denoising process progressively yields $x_{t-1}$ from $x_{t}$ in terms of $\epsilon_{t}=\mathcal{Z}_{\theta}(x_{t},t)$ until $t=0$. According to Bayes Theorem, the $x_{t-1}$ could be sampled through the following process, i.e.,
@@
where $\hat{x}_{0}$ is the estimated output from $x_{t}$ in terms of Eq.(1). To guide the $x_{t-1}$ towards the desired output $x_{0}$, the conditional guidance $y$ could be introduced through the following formula (Dhariwal & Nichol, 2021; Fei et al., 2023), i.e.,
@@
3.2 Framework Overview
To achieve OIR through PDM, the key lies in implementing $p_{\phi}(\cdot)$ to produce the appropriate guidances of gradients $\mathcal{G}$, guiding the generated image $\hat{x}_{0}$ towards the clean image $x_{0}$ of the test degraded image $y$. In this work, we implement $p_{\phi}(\cdot)$ through TDA and AIR, which are devised for open-set scenarios and image restoration, respectively. To overcome the challenges posed by open-set scenarios, TDA introduces a degradation-agnostic adapter during the test phase (i.e., denoising process of PDM) for adapting PDM to the unknown and unforeseen degradations. This adapter is implemented by a simple neural network $\phi$, and optimized once after each denoising step towards aligning the domain of the $\hat{x}_{0}$ to that of the $y$. Through the adapter, $\hat{x}_{0}$ could be translated to the corresponding degraded one $\phi(\hat{x}_{0})$, which could be further supervised towards the $y$, and in turn produce $\mathcal{G}$ to update the $\hat{x}_{0}$ towards the desired $x_{0}$. AIR is elaborated to conduct this supervised process for image restoration. Specifically, AIR is inspired by the observation that a PDM exhibits an intriguing temporal dynamic during the denoising process, i.e., $\hat{x}_{0}$ are generated from the unrecognizable contents to the perceptually rich contents, and finally the imperceptible details. Therefore, AIR dynamically adjusts the supervision strategies by gradually shifting the focuses from high- to low-level contents during the denoising process.
The framework of TAO is illustrated in Fig.2, where a PDM is fixed to progressively denoising the pure noise $x_{T}$ to the clean image $x_{0}$. For the test degraded image $y$, TAO first samples two maps of random noise at $t=T$, and thus two images will be generated after each denoising step. We refer to the image updated towards the clean image of the $y$ as the guided image $x^{g}_{t}$, and the other one as the auxiliary image $x^{a}_{t}$. We introduce the $x^{a}_{t}$ to prevent TDA from falling into trivial solutions, and assist AIR to guide the $x^{g}_{t}$ towards the desired clean image $x^{g}_{0}$. At each step $t$, the images $x^{g}_{t}$, $x^{a}_{t}$, and $y$ are first fed into TDA for optimizing the adapter $\phi$ through the objectives for domain alignment, and then sent into AIR for updating $x^{g}_{t}$ towards $x^{g}_{0}$, which best matches $y$ undergoing $\phi$, through the conditional guidance in Eq.(4).